# Yang-Baxter Hochschild Cohomology

### 01 Feb 2024 v2

@@ 5. Hopf algebra cohomology and braided algebra cohomology
Remark 5.1.
Suppose that $(\xi,\zeta)$ is a Hopf algebra $2$-cocycle. Then, if $\tilde{\mu}$ and $\tilde{\Delta}$ are the corresponding deformed structures, from the property of the antipode of a Hopf algebra, which holds true for the deformed structure as well, we have that $\tilde{\mu}(\mathbb{1}\otimes S)\tilde{\Delta}(x)=\tilde{\mu}(S\otimes\mathbb{1})\tilde{\Delta}=\xi(1)$, which in degree $1$ in $\hbar$ gives the equality
$\xi(x^{(1)}\otimes S(x^{(2)}))+\mu(\mathbb{1}\otimes S)\zeta(x)=\xi(S(x^{(1)})\otimes x^{(2)})+\mu(S\otimes\mathbb{1})\zeta(x)=0.$
Observe that in the case $\zeta=0$ this gives us that $\xi(x^{(1)}\otimes S(x^{(2)}))=\xi(S(x^{(1)})\otimes x^{(2)})=0$.
@@
$+\prescript{}{(1)}{y}\otimes S(\prescript{}{(2)}{y})x\prescript{}{(3)}{y}+\prescript{(1)}{}{y}\otimes S(\prescript{(2)}{}{y})x\prescript{(3)}{}{y}.$
Definition 5.2.
Let $H$ be a Hopf algebra with an antipode $S$. A 1-cochain $f\in C^{1}_{\rm Hf}(H,H)$ is called $S$-commuting if it satisfies $fS=Sf:H\rightarrow H$. The set of all $S$-commutative 1-cochains form a subgroup, called $S$-commutative 1-cochain group and is denoted by $C^{1}_{\rm HfS}(H,H)$. The quotient $Z^{2,i}_{\rm Hf}(H,H)/\delta^{1}_{\rm H}(C^{1}_{\rm HfS}(H,H))$ is denoted by $H^{2}_{\rm HfS}(H,H)$, and it is called the $S$-commuting Hopf algebra 2-cohomology group. Similar names and notations apply to boundary and cocycle groups.
Lemma 5.3.
If $\xi\in Z^{2,1}_{\rm Hf}(H,H)$ and $\zeta\in Z^{2,2}_{\rm H}(H,H)$ are cobounded by an $S$-commuting $1$-cochain, i.e. $\xi=\delta^{1,1}_{\rm Hf}(f)$ and $\zeta=\delta^{1,2}_{\rm Hf}(f)$ for some $f\in C^{1}_{\rm HfS}(H,H)$, where $\delta^{1,1}_{\rm Hf}\in{\rm Hom}(H^{\otimes 2},H)$ and $\delta^{1,2}_{\rm Hf}\in{\rm Hom}(H,H^{\otimes 2})$, then $(\Psi_{\zeta}(\xi),\xi)$ is cobounded by $f$, $\Psi_{\zeta}(\xi)=\delta^{1}_{\rm YB}(f)$.
@@
One computes
$\Psi_{\delta^{1,1}_{\rm Hf}(f)}(\delta^{1,2}_{\rm Hf}(f))(x\otimes y)$
$=$ $[y^{(1)}\otimes S(y^{(2)})f(x)y^{(3)}+y^{(1)}\otimes S(y^{(2)})xf(y^{(3)})-y^{(1)}S(y^{(2)})xf(y^{(3)})]$
$+[y^{(1)}\otimes f(S(y^{(2)}))xy^{(3)}+y^{(1)}S(y^{(2)})xf(y^{(3)})-y^{(1)}\otimes f(S(y^{(2)})y^{(3)})]$
$+[y^{(1)}\otimes S(f(y^{(2)})^{(1)})xf(y^{(2)})^{(2)}-y^{(1)}\otimes S(f(y^{(2)}))xy^{(3)}-y^{(1)}\otimes S(y^{(2)})xf(y^{(3)})]$
$+[f(y)^{(1)}\otimes S(f(y)^{(2)})xf(y^{(2)})^{(2)}-y^{(1)}\otimes S(f(y^{(2)})^{(1)})xf(y^{(3)})-f(y^{(1)})\otimes S(y^{(2)})xy^{(3)}]$
$=$ $\delta^{1}_{\rm YB}(f)(x\otimes y).$
The terms $y^{(1)}\otimes f(S(y^{(2)}))xy^{(3)}$ and $y^{(1)}\otimes S(f(y^{(2)})^{(1)})xf(y^{(3)})$ cancel due to the $S$-commuting condition on $f$. The remaining terms can be seen to constitute $\delta^{1}_{\rm YB}(\Psi_{\zeta}(\xi))$, completing the proof. ∎
We note that this does not imply that $\Psi$ provides a chain map, since $f$ is restricted to $S$-commuting 1-cochains.
Theorem 5.4.
Let $H$, $(\xi,\zeta)$ and $\Psi_{\zeta}$ be as above. Then, for each $\zeta$, $\Psi_{\zeta}$ induces a well defined morphism of cohomology groups $H^{2}_{\rm HfS}(H,H)\rightarrow H^{2}_{\rm YBH}(H,H)$ between $S$-commuting Hopf cohomology and YBH cohomology.
Proof.
Since $\xi\in H^{2}_{\rm Hf}(H,H)$ is mapped to $\psi\in H^{2}_{\rm YBH}(H,H)$, and
$\delta^{2,1}_{\rm Hf}(\xi)=\delta^{2}_{\rm H}(\psi),$
if $\xi$ is a 2-cocycle, the corresponding $\psi$ is also a 2-cocycle.
@@
We note that a $2$-cochain $\pi:H\otimes H\rightarrow H\otimes H$ is a YB $2$-cocycle for $R_{H}$ if and only if it is an infinitesimal deformation, i.e. if and only if $\tilde{R}=R_{H}+\hbar\pi$ is a YB operator for $H\otimes\mathbb{k}[[\hbar]]/(\hbar^{2})$. Let us consider the deformed structures $\tilde{\mu}=\mu+\hbar\xi$ and $\tilde{\Delta}=\Delta+\hbar\zeta$. They define a Hopf algebra structure (the infinitesimal deformation of $H$) since $(\xi,\zeta)$ is a Hopf $2$-cocycle by definition. Let $R_{\tilde{H}}$ be the YB operator defined through the adjoint map as in Section 2.3. Since $R_{\tilde{H}}$ can be written as $R_{\tilde{H}}=R_{H}+\hbar\Psi_{\zeta}(\xi)$ we have that $\Psi_{\zeta}(\xi)$ satisfies the YB $2$-cocycle condition, which shows that $(\Psi(\xi),\xi)$ is a braided algebra $2$-cocycle.
@@
We apply the same argument to show that $(\Psi_{\zeta}(\xi),\xi)$ satisfy the YI and IY $2$-cocycle conditions. By Lemma 3.6, the deformations $\tilde{\mu}=\mu+\hbar\psi=\mu+\hbar(\xi)$ and $\tilde{R}=R+\hbar\Psi_{\zeta}(\xi)$ define a braided algebra if and only if $(\Psi_{\zeta}(\xi),\xi)$ is a YI and IY 2-cocycle. By Lemma 2.6, if $\tilde{\mu}$ and $\tilde{\Delta}$ define a (infinitesimal) Hopf algebra, then $\mu$ and the adjoint YB operator $\tilde{R}$ define a braided algebra. Hence $(\Psi_{\zeta}(\xi),\xi)$ is a YI and IY 2-cocycle. ∎
Remark 5.5.
We observe that although the map in Theorem 5.4 relates $S$-commuting Hopf cohomology group and YBH cohomology, a substantially identical proof shows that the same correspondence maps nontrivial Hopf algebra $2$-cocycles to nontrivial YBH $2$-cocycles. However, this mapping might not descend to a well defined map between cohomology groups. This observation is useful in the examples below.
@@
By Theorem 3.8, if a cocycle deformation of a braided algebra by a pair of 2-cocycles $(\phi,\psi)$ defines a nontrivial deformation, then $(\phi,\psi)$ is a non-coboundary. While it is known that over fields of characteristic zero there are no nontrivial deformations for the group algebra [ChariPressley], for group algebras over more general rings there exist nontrivial deformations [SiegelWith]. Therefore, the corresponding braided algebras are deformed nontrivially. Hence in such cases the second cohomology group of YBH cohomology is nontrivial.
@@
Let us now consider a more general setting than that of Example 5.6 above. For an algebraic group $G$, let $\mathcal{F}(G)$ denote the algebra of regular functions. Then it is known (see [ChariPressley]) that there are nontrivial deformations of $\mathcal{F}(G)$ as a Hopf algebra, though the coalgebra structure has trivial deformation. Applying Remark 5.5 we obtain nontrivial deformations (and therefore nontrivial elements in the braided algebra second cohomology group according to Theorem 3.8) of the associated braided algebra.

### 10 Mar 2024 v3

@@ 5. Hopf algebra cohomology and braided algebra cohomology
Suppose that $(\xi,\zeta)$ is a Hopf algebra $2$-cocycle. It is known (see [ChariPressley]) that any bialgebra deformation of $H$ is equivalent to a deformation where unit and counit are not deformed. Thus, we will keep $\eta$ and $\epsilon$ to indicate unit and counit, respectively, without any deformation. This assumption is equivalent to that $(\xi,\zeta)$ is a normalized cocycle, i.e. a cocycle such that
$\xi(1\otimes x)=\xi(x\otimes 1)=0\quad{\rm and}\quad(\mathbb{1}\otimes\epsilon)\zeta(x)=(\epsilon\otimes\mathbb{1})\zeta(x)=0$
for all $x\in H$. Denote by $\hat{Z}_{\rm Hf}^{2}(H,H)$ the group of normalized Hopf algebra 2-cocycles.
Below we observe that for any normalized 2-cocycle $(\xi,\zeta)$, there is a unique deformation of the antipode $\tilde{S}=S+\hbar S^{\prime}$ that gives rise to an infinitesimal deformation of a Hopf algebra. The degree 1 term for the antipode condition is as follows:
(5) $\xi(\mathbb{1}\otimes S)\Delta+\mu(\mathbb{1}\otimes S)\zeta+\mu(\mathbb{1}\otimes S^{\prime})\Delta=\xi(S\otimes\mathbb{1})\Delta+\mu(S\otimes\mathbb{1})\zeta+\mu(S^{\prime}\otimes\mathbb{1})\Delta=0.$
We set $S^{\prime}(x)=-S(x^{(1)})\xi(x^{(2)}\otimes S(x^{(3)}))-S(x^{(1)})\mu((\mathbb{1}\otimes S)\zeta(x^{(2)}))$. Using this definition in Equation (5) we obtain
$\xi(\mathbb{1}\otimes S)\Delta(x)+\mu(\mathbb{1}\otimes S)\zeta(x)+\mu(\mathbb{1}\otimes S^{\prime})\Delta(x)$
$=$ $\xi(x^{(1)}\otimes S(x^{(2)}))+\mu(\mathbb{1}\otimes S)\zeta(x)-x^{(1)}S(x^{(2)})\xi(x^{(3)}\otimes S(x^{(4)}))$
$\hskip 156.49014pt-\mu(x^{(1)}\otimes S(x^{(2)})\mu((\mathbb{1}\otimes S)\otimes\zeta(x^{(3)}))$
$=$ $\xi(x^{(1)}\otimes S(x^{(2)}))+\mu(\mathbb{1}\otimes S)\zeta(x)-\epsilon(x^{(1)})\xi(x^{(2)}\otimes S(x^{(3)}))-\epsilon(x^{(1)})\mu((\mathbb{1}\otimes S)\zeta(x^{(2)})$
$=$ $0$
Applying the $2$-cocycle condition for $\xi$ on (the sum in Sweedler notation of terms) $S(x^{(1)}),x^{(2)},S(x^{(3)})$, and using the fact that $\xi$ is normalized, we find that
$\xi(S(x^{(1)})\otimes x^{(2)})S(x^{(3)})=S(x^{(1)})\xi(x^{(2)}\otimes S(x^{(3)})).$
Similarly, using the $2$-cocycle condition and normalization for $\zeta$ we also find that
$S(x^{(1)})\mu((\mathbb{1}\otimes S)\zeta(x^{(2)}))=\mu((S\otimes\mathbb{1})\zeta(x^{(1)}))S(x^{(2)}).$
Therefore, we can rewrite $S^{\prime}$ using these symmetries, and proceed in a way analogous to the previous result to obtain that $\xi(S\otimes\mathbb{1})\Delta+\mu(S\otimes\mathbb{1})\zeta+\mu(S^{\prime}\otimes\mathbb{1})\Delta=0$ as well, which completely shows that $S^{\prime}$ satisfies Equation (5). The fact that $S^{\prime}$ is unique, follows from uniqueness of antipodes in Hopf algebras. We will denote the normalized second cohomology group of $H$ with coefficients in $H$ with the symbol $\hat{H}^{2}_{\rm Hf}(H,H):=\hat{Z}_{\rm Hf}^{2}(H,H)/(B^{1}_{\rm Hf}(H,H)\cap\hat{Z}_{\rm Hf}^{2}(H,H))$.
@@
$+\prescript{}{(1)}{y}\otimes S(\prescript{}{(2)}{y})x\prescript{}{(3)}{y}+\prescript{(1)}{}{y}\otimes S(\prescript{(2)}{}{y})x\prescript{(3)}{}{y}+y^{(1)}\otimes S^{\prime}(y^{(2)})xy^{(3)}.$
Observe that since $S^{\prime}$ is uniquely associated to the pair $(\xi,\zeta)$, we do not indicate explicitly the dependence of $\Psi$ on $S^{\prime}$, but rather only express its dependence on $\xi$ and $\zeta$.
Theorem 5.1.
The map $\Psi$ defined above induces a well defined homomorphism $\Psi:\hat{H}^{2}_{\rm Hf}(H,H)\rightarrow H^{2}_{\rm YBH}(H,H)$, $(\xi,\zeta)\mapsto(\Psi_{\zeta}(\xi)),\xi)$. Moreover, if $[(\xi,\zeta)]\in H^{2}_{\rm Hf}(H,H)$ is such that $0\neq[\xi]\in H^{2}_{\rm H}(H,H)$, then $\Psi(\xi,\zeta)$ is the nonzero class in $H^{2}_{\rm YBH}(H,H)$.
@@
We note that a $2$-cochain $\phi:H\otimes H\rightarrow H\otimes H$ is a YB $2$-cocycle for $R_{H}$ if and only if it is an infinitesimal deformation, i.e. if and only if $\tilde{R}=R_{H}+\hbar\phi$ is a YB operator for $H\otimes\mathbb{k}[[\hbar]]/(\hbar^{2})$. Let us consider the deformed structures $\tilde{\mu}=\mu+\hbar\xi$ and $\tilde{\Delta}=\Delta+\hbar\zeta$. They define a Hopf algebra structure (the infinitesimal deformation of $H$) since $(\xi,\zeta)$ is a Hopf $2$-cocycle by definition. Let $R_{\tilde{H}}$ be the YB operator defined through the adjoint map as in Section 2.3 for the deformation $\tilde{H}$. Since $R_{\tilde{H}}$ can be written as $R_{\tilde{H}}=R_{H}+\hbar\Psi_{\zeta}(\xi)$ we have that $\Psi_{\zeta}(\xi)$ satisfies the YB $2$-cocycle condition. Since $\xi$ is a Hochschild $2$-cocycle by definition of Hopf cohomology, to show that $(\Psi(\xi),\xi)$ is a braided algebra $2$-cocycle, we need to show that the YI and IY $2$-cocycle conditions also hold.
@@
We apply the same argument as above to show that $(\Psi_{\zeta}(\xi),\xi)$ satisfy the YI and IY $2$-cocycle conditions. By Lemma 3.6, the deformations $\tilde{\mu}=\mu+\hbar\psi=\mu+\hbar(\xi)$ and $\tilde{R}=R+\hbar\Psi_{\zeta}(\xi)$ define a braided algebra if and only if $(\Psi_{\zeta}(\xi),\xi)$ is a YI and IY 2-cocycle. By Lemma 2.6, if $\tilde{\mu}$ and $\tilde{\Delta}$ define a (infinitesimal) Hopf algebra, then $\mu$ and the adjoint YB operator $\tilde{R}$ define a braided algebra. Hence $(\Psi_{\zeta}(\xi),\xi)$ is a YI and IY 2-cocycle.
We have therefore shown that $(\Psi(\xi),\xi)\in Z^{2}_{\rm YBH}(H,H)$. We now need to show that this map descends to cohomology. To do this, we need to prove that if $(\xi,\zeta)=(\delta^{1,1}_{\rm Hf}(f),\delta^{1,2}_{\rm Hf}(f))$ is cobounded and normalized, then $(\Psi_{\zeta}(\xi),\xi)$ is cobounded in YBH cohomology as well. Using the definition of Hochschild cohomology and YBH cohomology, we have that it is enough to show that $\Psi_{\zeta}(\xi)=\delta^{2}_{\rm YB}(g)$ for some $1$-cochain $g:H\rightarrow H$.
We compute first $\prescript{}{(1)}{x}\otimes\prescript{}{(2)}{x}\otimes\prescript{}{(3)}{x}$ and $\prescript{(1)}{}{x}\otimes\prescript{(2)}{}{x}\otimes\prescript{(3)}{}{x}$ for such a cobounded $\zeta$. We have
$\prescript{}{(1)}{x}\otimes\prescript{}{(2)}{x}\otimes\prescript{}{(3)}{x}$ $=$ $(\mathbb{1}\otimes\Delta)\delta_{\rm Hf}^{1,2}(f)(x)$
$=$ $-f(x)^{(1)}\otimes f(x)^{(2)}\otimes f(x)^{(3)}+f(x^{(1)})+f(x^{(1)})\otimes x^{(2)}\otimes x^{(3)}$
$+x^{(1)}\otimes f(x^{(2)})^{(1)}\otimes f(x^{(2)})^{(1)}\otimes f(x^{(2)})^{(2)}$
and
$\prescript{(1)}{}{x}\otimes\prescript{(2)}{}{x}\otimes\prescript{(3)}{}{x}$ $=$ $(\mathbb{1}\otimes\delta_{\rm Hf}^{1,2}(f))\Delta x$
$=$ $-x^{(1)}\otimes f(x^{(2)})^{(1)}\otimes f(x^{(2)})^{(2)}+x^{(1)}\otimes f(x^{(2)})\otimes x^{(3)}$
$+x^{(1)}\otimes x^{(2)}\otimes f(x^{(3)}).$
Before proving that $\Psi_{\zeta}(\xi)$ is a YB coboundary, we also show that for $(\xi,\zeta)=(\delta^{1,1}_{\rm Hf}(f),\delta^{1,2}_{\rm Hf}(f))$, we have $Sf-fS+S^{\prime}=0$, where $S^{\prime}$ is the degree 1 term of the unique antipode deformation associated with $(\xi,\zeta)$. In fact, we have
$S^{\prime}(x)$ $=$ $-S(x^{(1)})\xi(x^{(2)}\otimes S(x^{(3)}))-S(x^{(1)})\mu((\mathbb{1}\otimes S)\zeta(x^{(2)}))$
$=$ $f(S(x))-S(f(x))+\epsilon(f(x^{(2)}))S(x^{(1)})$
$=$ $f(S(x))-S(f(x)),$
where we have used the fact that $\epsilon(f(x^{(2)}))S(x^{(1)})=0$ because $\delta^{1}_{\rm Hf}(f)$ is normalized.
Now, we use these results to compute $\Psi_{\zeta}(\xi)$:
$\Psi_{\zeta}(\xi)(x\otimes y)$ $=$ $y^{(1)}\otimes S(y^{(2)})f(xy^{(2)})-y^{(1)}\otimes S(y^{(2)})xf(y^{(3)})-y^{(1)}\otimes S(y^{(2)})f(x)y^{(3)}$
$+y^{(1)}\otimes f(S(y^{(2)})xy^{(3)})-y^{(1)}\otimes S(y^{(2)})f(xy^{(3)})-y^{(1)}\otimes f(S(y^{(2)}))xy^{(3)}$
$-f(y)^{(1)}\otimes S(f(y)^{(2)})xf(y)^{(3)}+f(y^{(1)})\otimes S(y^{(2)})xy^{(3)}$
$+y^{(1)}\otimes S(f(y^{(2)})^{(1)})xf(y^{(2)})^{(2)}-y^{(1)}\otimes S(f(y^{(2)})^{(1)})xf(y^{(2)})^{(2)}$
$+y^{(1)}\otimes S(f(y^{(2)}))xy^{(3)}+y^{(1)}\otimes S(y^{(2)})xf(y^{(3)})+y^{(1)}\otimes S^{\prime}(y^{(2)})xy^{(3)}$
$=$ $-\delta_{\rm YBH}^{1}(f)(x\otimes y)-y^{(1)}\otimes f(S(y^{(2)}))xy^{(3)}$
$+y^{(1)}\otimes S(f(y^{(2)}))xy^{(3)}+y^{(1)}\otimes S^{\prime}(y^{(2)})xy^{(3)}$
$=$ $\delta_{\rm YB}^{1}(-f)(x\otimes y),$
where in the last step we have used the fact that $f(S(x))-Sf(x)+S^{\prime}(x)=0$. This shows that the map $\Psi$ induced by the correspondence $\Psi_{\zeta}(\xi)$ is well defined in cohomology.
To complete the proof of the theorem, we need to show that if $(\xi,\zeta)$ is a nontrivial deformation as an algebra, then its corresponding braided algebra deformation is nontrivial as well. This comes from the fact that if $(\Psi_{\zeta}(\xi),\xi)$ is cobounded in YBH cohomology, by definition it follows that $\xi$ is cobounded in Hochschild cohomology. ∎
@@
By Theorem 3.8, if a cocycle deformation of a braided algebra by a pair of 2-cocycles $(\phi,\psi)$ defines a nontrivial deformation, then $(\phi,\psi)$ is a non-coboundary. While it is known that over fields of characteristic zero there are no nontrivial deformations for the group algebra [ChariPressley], for group algebras over more general rings there exist nontrivial deformations of the algebra structure [SiegelWith]. Therefore, the corresponding braided algebras are deformed nontrivially. Hence in such cases the second cohomology group of YBH cohomology is nontrivial.
@@
Let us now consider a more general setting than that of Example 5.2 above. For an algebraic group $G$, let $\mathcal{F}(G)$ denote the algebra of regular functions. Then it is known (see [ChariPressley]) that there are nontrivial deformations of $\mathcal{F}(G)$ as a Hopf algebra, though the coalgebra structure has trivial deformation. Applying Theorem 5.1 we obtain nontrivial deformations (and therefore nontrivial elements in the braided algebra second cohomology group according to Theorem 3.8) of the associated braided algebra.